OpenStudy (anonymous):

[1.03] Simplify the following expression. 4x − 6y + 2z − 5x − 3y + 8z

OpenStudy (anonymous):

(4x-5x)(-6y-3y)(2z+8z)

OpenStudy (anonymous):

-1x-3y+10z

OpenStudy (anonymous):

lolhelpplz, you should explain how to do the problem instead of simply doing it and giving the answer.

OpenStudy (anonymous):

I gave them the second step.

OpenStudy (anonymous):

lol wrong

OpenStudy (anonymous):

Cara, I think you should try to do the problem on your own. Maybe you won't make the mistake that lol made.

OpenStudy (anonymous):

i think you should stfu and smoke some weed

OpenStudy (anonymous):

obviously i dont get it smartass

OpenStudy (anonymous):

Thaaat's cool. trolololo

OpenStudy (anonymous):

Cara, if you want to know how this is done, you need to group like terms. What are like terms? They have to have the same variable and the same power. 2x + 5x are like terms. 2x + 5y are not. 4y + y are like terms. 4y + \(y^2\) are not. So you look through your problem and you find like terms, and the group them. How do you group them? Just add 'em together. 2x + 5x = 7x 4y + y = 5y \(4x^2 - 8x^2 = -4x^2\)